OpenStudy (anonymous):

Describe the relationship among the numbers and symbols in two different ways. ( 24 + 36 ) /4

OpenStudy (anonymous):

@xMissAlyCatx

OpenStudy (xmissalycatx):

What do you mean by "describe the relationship"?

OpenStudy (anonymous):

umm... idk either. That is just what it says on this paper

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

does it say \[\frac{24+36}{4}\]?

OpenStudy (misty1212):

i am not exactly sure what you are being asked but it means "add 24 to 36, then divide the result by 4"

OpenStudy (misty1212):

i guess you could also write it as \[\frac{24}{4}+\frac{36}{4}\] which means divide 24 by 4, divide 36 by 4, then add

OpenStudy (anonymous):

oh... ok.... I guess but I am doing the expression, just so you know